# SalUn: Empowering Machine Unlearning via Gradient-based Weight Saliency in Both Image Classification and Generation

### 19 Feb 2024 v3

@@ 1 Introduction
In contrast to exact or certified MU, approximate unlearning has emerged as a more practical approach for ‘fast’ and ‘accurate’ unlearning. While the accuracy may not meet provable guarantees, it can b assessed using a broader range of practical metrics, such as membership inference attacks (Carlini et al., 2022), without necessitating data-model or algorithmic assumptions typically associated with certified unlearning. Despite the merits of practicality and efficiency, the performance of approximate unlearning can still exhibit significant variance. For example, influence unlearning (Izzo et al., 2021; Warnecke et al., 2021), built upon the influence function analysis of training data points (Koh & Liang, 2017), exhibits high performance variance due to the selection of hyperparameters required for influence function approximations, as well as the particular unlearning scenarios and evaluation metrics (Becker & Liebig, 2022), thereby raising concerns about instability in approximate unlearning methods. Other approximate unlearning methods, including Fisher forgetting (Golatkar et al., 2020), gradient ascent (Thudi et al., 2022a), and finetuning-based approaches (Warnecke et al., 2021; Jia et al., 2023), also face the similar challenge as will be illustrated later.
@@ 3 Preliminaries and Problem Statement
To be concrete, let $\mathcal{D}=\{\mathbf{z}_{i}\}_{i=1}^{N}$ denote the training dataset encompassing NN data points (including data feature $\mathbf{x}_{i}$ and label $\mathbf{y}_{i}$ for supervised learning). And let $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$ be the forgetting dataset. Its complement, denoted by $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$, is referred to as the remaining dataset. Prior to MU, we denote the original model as ${\bm{\theta}}_{\mathrm{o}}$}, trained on $\mathcal{D}$ using, e.g., empirical risk minimization (ERM). Consistent with existing literature (Thudi et al., 2022a; Jia et al., 2023), we regard Retrain as the MU’s gold standard, which involves training mode parameters ($\bm{\theta}$}) from scratch over $\mathcal{D}_{\mathrm{r}}$. Nonetheless, Retrain can be computationally demanding. Hence, the central challenge in MU is to acquire an unlearned model (referred to as ${\bm{\theta}}_{\mathrm{u}}$}) from ${\bm{\theta}}_{\mathrm{o}}$} on $\mathcal{D}_{\mathrm{f}}$ and/or $\mathcal{D_{\mathrm{r}}}$, so that it can serve as an accurate and computationally efficient substitute for Retrain. In what follows, we introduce two MU paradigms that are the primary focus of this work: MU for image classification and MU for image generation.
@@ 4 Challenges in Current Machine Unlearning Methods
The instability limitation. In evaluating the performance of MU methods, previous research has often assumed a fixed number of forgetting data, such as data points within an entire class or a fixed ratio of the training set. There has been limited evaluation exploring how the unlearning performance is affected by varying quantities of forgetting data. In Fig. 2(a), we investigate the unlearning performance gap relative to the gold standard Retrain, measured in terms of UA (unlearning accuracy), as a function of the quantity of forgetting data points. Note that a smaller performance gap is desirable. As we can see, the unlearning effectiveness of MU methods (①-⑤) observed at a 10% forgetting data quantity does not necessarily hold when the forgetting data quantity is increased to 50%. Similar instability in other performance metrics can also be observed in more experiment results later.
@@ 5 SalUn: Weight Saliency Is Possibly All You Need for MU
where $\mathbf{1}(\mathbf{g}\geq\gamma)$ is an element-wise indicator function which yields a value of 11 for the $i$-th element if $g_{i}\geq\gamma$ and 0 otherwise, $|\cdot|$ is an element-wise absolute value operation, and $\gamma>0$ is a hard threshold. In practice, we have observed that setting $\gamma$ to the median of the gradient vector $\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})\left.\right|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}}$ is a sufficiently effective choice. Based on (3), we express the unlearning model ${\bm{\theta}}_{\mathrm{u}}$} as
@@
Saliency-based unlearning (SalUn). Next, we introduce SalUn, which incorporates (4) into the unlearning process. One advantage of SalUn is its plug-and-play capability, allowing it to be applied on top of existing unlearning methods. In particular, we find that integrating weight saliency with the RL (random labeling) method provides a promising MU solution; See an ablation study in Table A1.
@@ 6.1 Experiment setups
Data, models, and unlearning setups. In image classification tasks, we focus on random data forgetting and evaluate its performance on the data-model (CIFAR-10, ResNet-18) (He et al., 2016; Krizhevsky et al., 2009). Additionally, we extend our evaluation to CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011), Tiny ImageNet (Le & Yang, 2015) datasets, and VGG-16 (Simonyan & Zisserman, 2014), Swin-T (Liu et al., 2021) architectures. We also consider the class-wise forgetting setup in image classification. See Appendix B.2 for these additional experiments.
@@ 6.2 Experiment results
Application to NSFW concept forgetting. Further, we assess the effectiveness of SalUn in concept-wise forgetting to eliminate the impact of NSFW (not safe for work) concepts introduced through inappropriate image prompts (I2P) (Schramowski et al., 2023). Specifically, we generate images from the open-source SD V1.4 using prompts provided by I2P and classify them into various corresponding nude body parts using the NudeNet detector (Bedapudi, 2019). Our goal is to leverage an MU method to effectively erase the influence of the nudity-related prompts in SD. Fig. 5 presents the unlearning performance of different unlearning methods, including SalUn, and the ESD and FMT baselines introduced in Table 2. Here, the unlearning effectiveness is measured by the amount of nudity-related image generations using the unlearned SD with I2P prompts. We also present the performance of the original SD model for comparison. CFDeleted As we can see, SalUn generates the fewest harmful images across all the nude body part classes. In particular, it significantly outperforms ESD (the second best-performing method) in ‘male breast’ and ‘female breast’. Additionally, in the absence of unlearning, the original SD V1.4 can generate a large number of harmful images, underscoring the importance of MU for image generation. Fig. 6 provides a set of generated images using I2P prompts, illustrating the generation differences among the original SD model and its various unlearned versions.
@@ Appendix A Soft-thresholding SalUn
where $i=1$ (or 22) corresponds to the objective function of SalUn for classification (6) (or generation (7)), and $\beta>0$ is a regularization parameter. Note that unlike (6) and (7), the objective function of SalUn is defined over the entire model weights $\bm{\theta}$} since there is no weight saliency map known in advance, i.e., letting $\mathbf{m}_{\mathrm{S}}=\mathbf{1}$ in (4). Problem (8) can be efficiently solved using the proximal gradient algorithm (Parikh et al., 2014), wherein the $\ell_{1}$ penalty term is handled efficiently through a closed-form proximal operation known as soft-thresholding. In contrast to the hard-thresholding implementation of SalUn, the soft-thresholding approach requires tuning an additional hyperparameter $\beta$, with the trade-off of not resorting to the hard threshold $\gamma$ in (3). In practice, we find that employing a linearly decaying scheduler for $\beta$ (i.e., promoting weight saliency sparsity during the early unlearning stages) yields better results than other schemes, approaching the performance of SalUn based on hard thresholding.
@@ B.1 Additional training and unlearning settings
For DDPM, the unlearning settings are as follows: For Retrain, it undergoes training for 80,000 iterations with Adam and a learning rate of $10^{-4}$. The batch size is set to 128. In the case of SalUn, it is trained for 1,000 iterations with a learning rate of $10^{-4}$, $\alpha$ set to $10^{-3}$, and a batch size of 128. The sparsity of weight saliency is maintained at 50%. The sampling settings include 1,000 time steps and a conditional scale of 2.0.
@@
For NSFW removal, we initially employ SD V1.4 to generate 800 images as $\mathcal{D}_{\mathrm{f}}$ using the prompt ‘a photo of a nude person’ and an additional 800 images as $\mathcal{D}_{\mathrm{r}}$ using the prompt ‘a photo of a person wearing clothes.’ Throughout the Unlearning process, we utilize ‘a photo of a nude person’ to derive the weight saliency mask for NSFW concept. Subsequently, we regard ‘a photo of a nude person’ as the concept to be forgotten and make correction using the concept ‘a photo of a person wearing clothes.’
@@
Additionally, Fig. A2 shows the examples of generated images when forgetting the class ‘airplane’. It’s evident that opting for a 10% saliency sparsity can effectively forget the class ‘airplane,’ but it leads to a decrease in image generation quality. On the other hand, using a 90% saliency sparsity fails to forget the class. In Fig. A3, we present how the ‘optimal’ sparsity level can change vs. the number of forgetting data points, focusing on the case of MU for image classification. The optimal sparsity is determined through a grid search with a sparsity interval of 10%. It’s important to note that the choice of a 50% sparsity ratio is not universal, and when dealing with a larger amount of forgetting data, a lower sparsity ratio is preferred. This is because a larger forgetting data set requires more weight updates, leading to a lower saliency sparsity.
@@ B.2 Additional classification results
Performance of Tiny ImageNet dataset
We conducted additional experiments on the Tiny ImageNet dataset (Le & Yang, 2015) with a higher resolution ($64\times 64$) than CIFAR-10 and CIFAR-100 in Table A4. We focused on evaluating our method against the baseline $\ell_{1}$-sparse and Retrain. Compared to $\ell_{1}$-sparse, SalUn demonstrates smaller gaps in terms of UA and MIA, with comparable RA and TA gaps. This shows an improved unlearning efficacy over $\ell_{1}$-sparse while preserving the model’s generalization ability post unlearning.

### 01 Mar 2024 v4

@@ 1 Introduction
In contrast to exact or certified MU, approximate unlearning has emerged as a more practical approach for ‘fast’ and ‘accurate’ unlearning. While the accuracy may not meet provable guarantees, it can b assessed using a broader range of practical metrics, such as membership inference attacks (Carlini et al., 2022), without necessitating data-model or algorithmic assumptions typically associated with certified unlearning. Despite the merits of practicality and efficiency, the performance of approximate unlearning can still exhibit significant variance. For example, influence unlearning (Izzo et al., 2021; Warnecke et al., 2021), built upon the influence function analysis of training data points (Koh & Liang, 2017), exhibits high-performance variance due to the selection of hyperparameters required for influence function approximations, as well as the particular unlearning scenarios and evaluation metrics (Becker & Liebig, 2022), thereby raising concerns about instability in approximate unlearning methods. Other approximate unlearning methods, including Fisher forgetting (Golatkar et al., 2020), gradient ascent (Thudi et al., 2022a), and finetuning-based approaches (Warnecke et al., 2021; Jia et al., 2023), also face the similar challenge as will be illustrated later.
@@ 3 Preliminaries and Problem Statement
To be concrete, let $\mathcal{D}=\{\mathbf{z}_{i}\}_{i=1}^{N}$ denote the training dataset encompassing $N$ data points (including data feature $\mathbf{x}_{i}$ and label $\mathbf{y}_{i}$ for supervised learning). And let $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$ be the forgetting dataset. Its complement, denoted by $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$, is referred to as the remaining dataset. Prior to MU, we denote the original model as ${\bm{\theta}}_{\mathrm{o}}$}, trained on $\mathcal{D}$ using, e.g., empirical risk minimization (ERM). Consistent with existing literature (Thudi et al., 2022a; Jia et al., 2023), we regard Retrain as the MU’s gold standard, which involves training mode parameters ($\bm{\theta}$}) from scratch over $\mathcal{D}_{\mathrm{r}}$. Nonetheless, Retrain can be computationally demanding. Hence, the central challenge in MU is to acquire an unlearned model (referred to as ${\bm{\theta}}_{\mathrm{u}}$}) from ${\bm{\theta}}_{\mathrm{o}}$} on $\mathcal{D}_{\mathrm{f}}$ and/or $\mathcal{D_{\mathrm{r}}}$, so that it can serve as an accurate and computationally efficient substitute for Retrain. In what follows, we introduce two MU paradigms that are the primary focus of this work: MU for image classification and MU for image generation.
@@ 4 Challenges in Current Machine Unlearning Methods
The instability limitation. In evaluating the performance of MU methods, previous research has often assumed a fixed number of forgetting data, such as data points within an entire class or a fixed ratio of the training set. There has been limited evaluation exploring how the unlearning performance is affected by varying quantities of forgetting data. In Fig. 2(a), we investigate the unlearning performance gap relative to the gold standard Retrain, measured in terms of UA (unlearning accuracy), as a function of the quantity of forgetting data points. Note that a smaller performance gap is desirable. As we can see, the unlearning effectiveness of MU methods (①-⑤) observed at a 10% forgetting data quantity does not necessarily hold when the forgetting data quantity is increased to 50%. Similarly, the instability can also be observed in other performance metrics and will show in experiment results later.
@@ 5 SalUn: Weight Saliency Is Possibly All You Need for MU
where $\mathbf{1}(\mathbf{g}\geq\gamma)$ is an element-wise indicator function which yields a value of $1$ for the $i$-th element if $g_{i}\geq\gamma$ and 0 otherwise, $|\cdot|$ is an element-wise absolute value operation, and $\gamma>0$ is a hard threshold. In practice, we have observed that setting $\gamma$ to the median of the gradient vector $\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})\left.\right|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}}$ is a sufficiently effective choice. Based on (3), we express the unlearning model ${\bm{\theta}}_{\mathrm{u}}$} as
@@
Saliency-based unlearning (SalUn). Next, we introduce SalUn, which incorporates (4) into the unlearning process. One advantage of SalUn is its plug-and-play capability, allowing it to be applied on top of existing unlearning methods. In particular, we find that integrating weight saliency with the RL (random labeling) method provides a promising MU solution; See the ablation study in Table A1.
@@ 6.1 Experiment setups
Data, models, and unlearning setups. In image classification tasks, we focus on random data forgetting and evaluate its performance on the data model (CIFAR-10, ResNet-18) (He et al., 2016; Krizhevsky et al., 2009). Additionally, we extend our evaluation to CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011), Tiny ImageNet (Le & Yang, 2015) datasets, and VGG-16 (Simonyan & Zisserman, 2014), Swin-T (Liu et al., 2021) architectures. We also consider the class-wise forgetting setup in image classification. See Appendix B.2 for these additional experiments.
@@ 6.2 Experiment results
Application to NSFW concept forgetting. Further, we assess the effectiveness of SalUn in concept-wise forgetting to eliminate the impact of NSFW (not safe for work) concepts introduced through inappropriate image prompts (I2P) (Schramowski et al., 2023). Specifically, we generate images from the open-source SD V1.4 using prompts provided by I2P and classify them into various corresponding nude body parts using the NudeNet detector (Bedapudi, 2019). Our goal is to leverage an MU method to effectively erase the influence of the nudity-related prompts in SD. Fig. 5 presents the unlearning performance of different unlearning methods, including SalUn, and the ESD and FMT baselines introduced in Table 2. Here, the unlearning effectiveness is measured by the amount of nudity-related image generations using the unlearned SD with I2P prompts. We also present the performance of the original SD model for comparison. SalUn generates the fewest harmful images across all the nude body part classes. In particular, it significantly outperforms ESD (the second best-performing method) in ‘male breast’ and ‘female breast’. Additionally, in the absence of unlearning, the original SD V1.4 can generate a large number of harmful images, underscoring the importance of MU for image generation. Fig. 6 provides a set of generated images using I2P prompts, illustrating the generation differences among the original SD model and its various unlearned versions.
@@ Appendix A Soft-thresholding SalUn
where $i=1$ (or $2$) corresponds to the objective function of SalUn for classification (6) (or generation (7)), and $\beta>0$ is a regularization parameter. Note that unlike (6) and (7), the objective function of SalUn is defined over the entire model weights $\bm{\theta}$} since there is no weight saliency map known in advance, i.e., letting $\mathbf{m}_{\mathrm{S}}=\mathbf{1}$ in (4). Problem (8) can be efficiently solved using the proximal gradient algorithm (Parikh et al., 2014), wherein the $\ell_{1}$ penalty term is handled efficiently through a closed-form proximal operation known as soft-thresholding. In contrast to the hard-thresholding implementation of SalUn, the soft-thresholding approach requires tuning an additional hyperparameter $\beta$, with the trade-off of not resorting to the hard threshold $\gamma$ in (3). In practice, we find that employing a linearly decaying scheduler for $\beta$ (i.e., promoting weight saliency sparsity during the early unlearning stages) yields better results than other schemes, approaching the performance of SalUn based on hard thresholding.
@@ B.1 Additional training and unlearning settings
For DDPM, the unlearning settings are as follows: For Retrain, it undergoes training for 80,000 iterations with Adam and a learning rate of $10^{-4}$. The batch size is set to 128. In the case of SalUn, it is trained for 1,000 iterations with a learning rate of $10^{-4}$, $\alpha$ set to $10^{-3}$ and a batch size of 128. The sparsity of weight saliency is maintained at 50%. The sampling settings include 1,000 time steps and a conditional scale of 2.0.
@@
For NSFW removal, we initially employ SD V1.4 to generate 800 images as $\mathcal{D}_{\mathrm{f}}$ using the prompt ‘a photo of a nude person’ and an additional 800 images as $\mathcal{D}_{\mathrm{r}}$ using the prompt ‘a photo of a person wearing clothes.’ Throughout the Unlearning process, we utilize ‘a photo of a nude person’ to derive the weight saliency mask for the NSFW concept. Subsequently, we regard ‘a photo of a nude person’ as a concept to be forgotten and make corrections using the concept ‘a photo of a person wearing clothes.’
@@
Additionally, Fig. A2 shows the examples of generated images when forgetting the class ‘airplane’. Opting for a 10% saliency sparsity can effectively forget the class ‘airplane,’ but it leads to a decrease in image generation quality. On the other hand, using a 90% saliency sparsity fails to forget the class. In Fig. A3, we present how the ‘optimal’ sparsity level can change vs. the number of forgetting data points, focusing on the case of MU for image classification. The optimal sparsity is determined through a grid search with a sparsity interval of 10%. It’s important to note that the choice of a 50% sparsity ratio is not universal, and when dealing with a larger amount of forgetting data, a lower sparsity ratio is preferred. This is because a larger forgetting data set requires more weight updates, leading to a lower saliency sparsity.
@@ B.2 Additional classification results
Performance of Tiny ImageNet dataset.
We conducted additional experiments on the Tiny ImageNet dataset (Le & Yang, 2015) with a higher resolution ($64\times 64$) than CIFAR-10 and CIFAR-100 in Table A4. We focused on evaluating our method against the baseline $\ell_{1}$-sparse and Retrain. Compared to $\ell_{1}$-sparse, SalUn demonstrates smaller gaps in terms of UA and MIA, with comparable RA and TA gaps. This shows an improved unlearning efficacy over $\ell_{1}$-sparse while preserving the model’s generalization ability post-unlearning.